# Resource theories with completely passive states and passive states as free states

### v3

@@ Abstract
Work extractable from quantum system is a resource that can be related to some resource theory. By considering completely passive states and passive states as free states, we formulate resource theories where the maximum work extractable is a monotone, showing how the definition of a temperature plays a pivotal role in the resource theories.
@@ I Introduction
Work extraction processes are of fundamental importance in quantum thermodynamics [1], for instance recently they played a certain role in the context of quantum batteries [2]. For finite thermally isolated quantum systems, the maximum work extractable through a cycle is upper bounded by a free energy that is obtained by minimizing the final energy with the constraint that the von Neumann entropy remains constant [3]. In general, the maximum work is equal to the quantum ergotropy [4] and the states with zero ergotropy are passive. The free energy can be extracted in the limit of an infinite number of copies [5] and the states that are passive for any number of copies are completely passive. In this paper, the maximum work extractable is viewed as a resource, and thus linked to the paradigm of the resource theories (see, e.g., Ref. [6] for the role of resource theories in quantum thermodynamics). For instance, a paradigmatic resource theory views the entanglement as a resource [7]. In this resource theory, states without entaglement are separable states, which are viewed as free states. Furthermore, local operations and classical communication do not generate entanglement, leaving separable states separable, and thus are viewed as free operations. Similarly, other resource theories can be formulated, for instance a resource theory of quantum coherence [8]. Furthermore, a resource theory of quantum thermodynamics has been introduced in Ref. [9], where the free operations are thermal operations and do not increase the nonequilibrium Helmholtz free energy (see, e.g., Ref. [10]). From a microscopic point of view, work can be extracted through thermal operations only when the system interacts with an environment at a certain initial temperature. In contrast, we focus on reversible work extractions, discussing two resource theories where the resources are the free energy and the ergotropy, respectively.
Thus, after introducing some preliminary notions in Sec. II, we show how the free energy and the ergotropy can be viewed as distance-based monotone of the corresponding resource theories in Secs. III and IV, respectively. A unified picture follows for the two resource theories by expressing the monotones in terms of a unique distance. In particular, the distance is achieved by multiplying the relative entropy by a nonequilibrium temperature. Then, free operations can increase the nonequilibrium temperature of the free states at most by a certain factor. We examine the monotonicity properties of the free energy and the ergotropy, the contractivity of the distance and other properties, also with the help of specific free operations. Finally, we summarize and discuss further the results achieved in Sec. V.
@@ II.1 Resource theories
where $D(\rho,\sigma)$ is a distance, i.e., a nonnegative functional such that $D(\rho,\sigma)=0$ if and only if $\rho=\sigma$, so that the monotone satisfies (i). For instance, the distance can be chosen to be the quantum relative entropy $S(\rho||\sigma)=\text{Tr}\left\{\rho(\ln\rho-\ln\sigma)\right\}$, so that it is contractive, $D(\Lambda(\rho),\Lambda(\sigma))\leq D(\rho,\sigma)$ for any CPTP map, and the monotone satisfies (ii).
@@
We can formulate different resource theories depending on what the resource is. For instance, the quantum entanglement can be viewed as a resource, such that the free state are separable states and the free operations are obtained through local operations and classical communication [7]. When the quantum coherence is viewed as a resource, the incoherent states, i.e., the states diagonal with respect to a fixed basis, are free states, and the largest set of free operations is formed by the maximally incoherent operations, which are defined such that they map incoherent states into incoherent states [8]. In the resource theory of quantum thermodynamics [10], the free operations are the so-called thermal operations, which are defined as $\Lambda_{th}(\rho)=\text{Tr}_{E}\left\{U\rho\otimes\gamma^{E}_{\beta}U^{\dagger}\right\}$ with unitary operators $U$ that commutate with the total Hamiltonian. In detail, the thermal operations act on the Hilbert space of a system $S$ having Hamiltonian $H$, and are defined by performing a Stinespring dilation with an environment $E$ having Hamiltonian $H_{E}$ and prepared in the Gibbs state $\gamma^{E}_{\beta}=e^{-\beta H_{E}}/Z_{E}$ at the temperature $T=\beta^{-1}$, with $Z_{E}=\text{Tr}\left\{e^{-\beta H_{E}}\right\}$, and with a global unitary operator $U$ such that $[U,H+H_{E}]=0$. The thermal operations preserve the Gibbs state of the system $S$, $\Lambda_{th}(\gamma_{\beta})=\gamma_{\beta}$, thus the Gibbs state $\gamma_{\beta}=e^{-\beta H}/Z$, with $Z=\text{Tr}\left\{e^{-\beta H}\right\}$, is the free state. It results that not all the CPTP maps that preserve the Gibbs state $\gamma_{\beta}$, forming the set of Gibbs-preserving operations, are thermal operations (see, e.g., Ref. [11]). Furthermore, the nonequilibrium Helmholtz free energy defined as $F_{\beta}(\rho)=TS(\rho||\gamma_{\beta})-T\ln Z$ does not increase under the action of the thermal operations, i.e., $F_{\beta}(\Lambda_{th}(\rho))\leq F_{\beta}(\rho)$, so that it is a monotone. The maximum work extractable is given by the difference $\Delta F_{\beta}(\rho)=F_{\beta}(\rho)-F_{\beta}(\gamma_{\beta})=TS(\rho||\gamma_{\beta})$. We note that in this case a work extraction is microscopically performed through an environment $E$, which initially is in the equilibrium state $\gamma^{E}_{\beta}$. In contrast, in the following we focus on a reversible work extraction, such that the system $S$ is thermally isolated. Thus, in order to extract a nonzero average work, the unitary time-evolution operator must not commutate with the Hamiltonian. We aim to define some free operations that do not increase the maximum work extractable, which is viewed as a resource. In particular, we find that in the definition of these free operations some sort of temperature plays a pivotal role.
@@ II.2 Work extraction
Following Ref. [4], we consider a finite quantum system $S$ having a Hilbert space $\mathcal{H}$ with dimension $d$. The Hamiltonian is expressed as
@@
with $\gamma_{\rho}=\gamma_{\beta(\rho)}$ and $\beta(\rho)$ solution of the nonlinear equation for $\beta$, $S(\rho_{\beta})=S(\rho)$. Here, we simply refer to $F(\rho)$ as free energy, as in Ref. [3]. In general, the maximum work extractable is achieved by maximizing Eq. (4) over all the unitary cycles and it is equal to the ergotropy [4], here denoted with $\mathcal{E}(\rho)$, i.e.,
@@
where $P_{\rho}=\sum_{k}r_{k}|\epsilon_{k}\rangle\langle\epsilon_{k}|$ is a passive state. In particular, the maximum average work $\mathcal{E}(\rho)$ is obtained for a unitary operator $U_{\rho}=\sum_{k}e^{i\phi_{k}}|\epsilon_{k}\rangle\langle r_{k}|$, where $\phi_{k}$ are some arbitrary real phases, so that the final state $\rho^{\prime}=U_{\rho}\rho U_{\rho}^{\dagger}$ is the passive state $P_{\rho}$. In particular, we note that $P_{U\rho U^{\dagger}}=P_{\rho}$ for any unitary operator $U$. In general, the set $\Pi$ of the passive states is formed by all the states $P$ having zero ergotropy, i.e., such that $\mathcal{E}(P)=0$. In other words, the ergotropy $\mathcal{E}(\rho)$ is zero if and only if the state $\rho$ is passive. Passive states are Gibbs states $\gamma_{\beta}$ with $\beta\geq 0$, and in general the set $\Pi$ is formed by all the states $P=\sum_{k}p_{k}|\epsilon_{k}\rangle\langle\epsilon_{k}|$ with $p_{k}\geq p_{k+1}$.
Let us consider an ensemble of $n$ copies of the system $S$, as in Ref. [5], where any copy has the Hamiltonian $H$ and is prepared in the state $\rho$. The state $\rho$ is completely passive if the state $\rho^{\otimes n}$ is passive with respect to the total Hamiltonian for any $n$. In general, the set $C$ of the completely passive states is formed by all the Gibbs states $\gamma_{\beta}$ with $\beta\geq 0$. This can be understood by noting that $\mathcal{E}(\rho^{\otimes n})\leq nF(\rho)$ for any $n$. In particular, it is possible to perform a global unitary cycle such that [5, 12]
@@
as $n\to\infty$. In the end, we note that, since $\gamma_{U\rho U^{\dagger}}=\gamma_{\rho}$ for any unitary operator $U$, the free energy and the ergotropy are related by the equation
@@ III Completely passive states as free states
Then, $\mathcal{O}_{cp}\subseteq\mathcal{M}_{cp}=\{\Lambda|\Lambda(C)\subseteq C\}$ and $M_{cp}(\rho)=\min_{\gamma\in C}S(\rho||\gamma)$ is a monotone, since $M_{cp}(\Lambda(\rho))\leq M_{cp}(\rho)$ for all $\Lambda\in\mathcal{M}_{cp}$. In contrast, by defining the temperature functional $T(\gamma)$ such that $T(\gamma_{\beta})=\beta^{-1}$, the free energy $F(\rho)$ can be expressed as [3]
@@
where we defined the contraction factor
@@
Thus, the set of free operations $\mathcal{O}_{cp}$ is formed by all the CPTP maps $\Lambda$ that satisfy the conditions (F,i) and (F,iii), where (F,iii) is the condition in Eq. (12). We can see that the condition (F,i) does not imply (F,ii) by considering a map $\Lambda_{\beta}$ such that $\Lambda_{\beta}(\gamma)=\gamma_{\beta}$. Similarly to the maps considered in Refs. [11, 13], the map $\Lambda_{\beta}$ can be realized as
@@
for all $\gamma\in C$, where $\sigma_{i}=K_{i}\rho K_{i}^{\dagger}/q_{i}$, $q_{i}=\text{Tr}\left\{K^{\dagger}_{i}K_{i}\rho\right\}$, $\gamma_{i}=K_{i}\gamma K_{i}^{\dagger}/\text{Tr}\left\{K^{\dagger}_{i}K_{i}\gamma\right\}$ and $\Lambda(\rho)=\sum_{i}K_{i}\rho K^{\dagger}_{i}=\sum_{i}q_{i}\sigma_{i}$. For instance, Eq. (25) is satisfied if $T(\gamma_{i})\leq T(\gamma)$ for all $i$, since in general $S(\rho||\gamma)\geq\sum_{i}q_{i}S(\sigma_{i}||\gamma_{i})$ (see, e.g., Ref. [14]).
@@
If we consider two systems with Hamiltonian $H_{1}$ and $H_{2}$, and the composite system with Hamiltonian $H=H_{1}+H_{2}$, we recall that the free energy is super-additive under tensor products [3], i.e., $F(\rho_{1}\otimes\rho_{2})\geq F(\rho_{1})+F(\rho_{2})$, and it is additive if $\beta(\rho_{1})=\beta(\rho_{2})$, thus it cannot be considered as a measure. Furthermore, the set $C$ is not convex for $d>2$, i.e., the convex combination of two Gibbs states is not always a Gibbs state, and the free energy $F(\rho)$ is not convex, e.g., for $\rho=\sum_{i}p_{i}\gamma_{i}\notin C$, we get $F(\rho)>\sum_{i}p_{i}F(\gamma_{i})=0$ if $\gamma_{i}\in C$ for all $i$.
@@ IV Passive states as free states
Let us try to extend the results achieved to the case of free states that are passive states. We consider the set $\Pi$ of the passive states, $\Pi=\{P=\sum_{k}p_{k}|\epsilon_{k}\rangle\langle\epsilon_{k}||p_{k}\geq p_{k+1}\}$, as the set of the free states. The set $\mathcal{O}_{p}$ of the free operations is formed by all the CPTP maps $\Lambda$ that satisfy (E,i) $\Lambda(\Pi)\subseteq\Pi$ and (E,ii) $\mathcal{E}(\Lambda(\rho))\leq\mathcal{E}(\rho)$ for all $\rho$. Similarly to the free energy case, the condition (E,ii) implies (E,i). Then, $\mathcal{O}_{p}\subseteq\mathcal{M}_{p}=\{\Lambda|\Lambda(\Pi)\subseteq\Pi\}$ and $M_{p}(\rho)=\min_{P\in\Pi}S(\rho||P)$ is a monotone, since $M_{p}(\Lambda(\rho))\leq M_{p}(\rho)$ for all $\Lambda\in\mathcal{M}_{p}$. We note that the ergotropy is equal to the free energy, $\mathcal{E}(\rho)=F(\rho)$, when $P_{\rho}$ is a completely passive state, i.e., $P_{\rho}=\gamma_{\rho}$. Then, to generalize the results achieved for the free energy to the case of the passive states as free states, we try to define a nonequilibrium temperature for the passive state $P$ such that it is equal to the temperature functional $T(P)$ when $P=\gamma\in C$. Thus, from the monotone $M_{p}(\rho)$, similarly to Eq. (11), we define a functional
@@
where $T(P|\rho)$ is some nonequilibrium temperature for the passive state $P$. Of course there are infinite ways to define the nonequilibrium temperature of a non-Gibbs passive state $P$, such that $T(P)=\beta^{-1}$ when $P=\gamma_{\beta}$. When we perform the minimization in Eq. (29), we can select a certain definition of the nonequilibrium temperature, among all the possible definitions, depending on the state $\rho$, achieving in this way a nonequilibrium temperature $T(P|\rho)$ that depends on $\rho$. The minimum in Eq. (29) can be achieved for a unique passive state $P^{*}$, such that $\Phi(\rho)=T(P^{*}|\rho)S(\rho||P^{*})$. Given a state $\rho$, we can choose the nonequilibrium temperature $T(P|\rho)$ such that $P^{*}=P_{\rho}$ and the functional in Eq. (29) is equal to the ergotropy, i.e., $\Phi(\rho)=\mathcal{E}(\rho)$. In detail, we find that $P^{*}=P_{\rho}$ and
@@
where we defined the contraction factor

### v4

@@ Abstract
Work extractable from quantum system can be viewed as a resource to perform some tasks. Here, we try to related the work extractable to some resource theory. To do this, we consider completely passive states and passive states as free states and we formulate resource theories where the maximum work extractable is a monotone, by identifying the free operations. In particular, we study how the free operations act on the free states, showing how a sort of temperature plays a pivotal role in these resource theories.
@@ I Introduction
Work extraction processes are of fundamental importance in quantum thermodynamics bookthermo , for instance recently they played a certain role in the context of quantum batteries Campaioli23 . For finite thermally isolated quantum systems, the maximum work extractable through a cycle is upper bounded by a free energy that is obtained by minimizing the final energy with the constraint that the von Neumann entropy remains constant Bera19 . In general, the maximum work is equal to the quantum ergotropy Allahverdyan04 and the states with zero ergotropy are passive. The free energy can be extracted in the limit of an infinite number of copies Alicki13 and the states that are passive for any number of copies are completely passive. In this paper, the maximum work extractable is viewed as a resource, and thus linked to the paradigm of the resource theories (see, e.g., Ref. Goold16 for the role of resource theories in quantum thermodynamics). For instance, a paradigmatic resource theory views the entanglement as a resource Horodecki09 . In this resource theory, states without entaglement are separable states, which are viewed as free states. Local operations and classical communication do not generate entanglement, leaving separable states separable, and thus are viewed as free operations. Similarly, other resource theories can be formulated, for instance a resource theory of quantum coherence Streltsov17 . Furthermore, a resource theory of quantum thermodynamics has been introduced in Ref. Janzing00 , where the free operations are thermal operations and do not increase the nonequilibrium Helmholtz free energy (see, e.g., Ref. Lostaglio19 ). From a microscopic point of view, work can be extracted through thermal operations only when the system interacts with an environment at a certain initial temperature. In contrast, we focus on reversible work extractions, discussing two resource theories where the resources are the free energy and the ergotropy, respectively. Our main motivation is to understand what is the general structure of the processes that can be performed without generating free energy or ergotropy so as to use these resources.
Thus, after introducing some preliminary notions in Sec. II, we show how the free energy and the ergotropy can be viewed as distance-based monotone of the corresponding resource theories in Secs. III and IV, respectively. A unified picture follows for the two resource theories by expressing the monotones in terms of a unique distance (in broad sense). In particular, the distance is achieved by multiplying the relative entropy by a sort of nonequilibrium temperature. Then, free operations can increase the nonequilibrium temperature of the free states at most by a certain factor. We examine the monotonicity properties of the free energy and the ergotropy, the contractivity of the distance and other properties, also with the help of specific free operations. We also introduce some other monotones for the resource theories in Sec. V. Finally, we summarize and discuss further the results achieved in Sec. VI.
@@ II.1 Resource theories
where $D(\rho,\sigma)$ is a distance. Here “distance” is understood in a rather broad term, i.e., $D(\rho,\sigma)$ is a nonnegative functional such that $D(\rho,\sigma)=0$ if and only if $\rho=\sigma$, so that the monotone satisfies (i). For instance, the distance can be chosen to be the quantum relative entropy $S(\rho||\sigma)=\text{Tr}\left\{\rho(\ln\rho-\ln\sigma)\right\}$, so that it is contractive, $D(\Lambda(\rho),\Lambda(\sigma))\leq D(\rho,\sigma)$ for any CPTP map, and the monotone satisfies (ii).
@@
We can formulate different resource theories depending on what the resource is. For instance, the quantum entanglement can be viewed as a resource, such that the free state are separable states and the free operations are obtained through local operations and classical communication Horodecki09 . When the quantum coherence is viewed as a resource, the incoherent states, i.e., the states diagonal with respect to a fixed basis, are free states, and the largest set of free operations is formed by the maximally incoherent operations, which are defined such that they map incoherent states into incoherent states Streltsov17 . In particular, we are going to follow this approach in the next sections. In the resource theory of quantum thermodynamics Lostaglio19 , the free operations are the so-called thermal operations, which are defined as $\Lambda_{th}(\rho)=\text{Tr}_{E}\left\{U\rho\otimes\gamma^{E}_{\beta}U^{\dagger}\right\}$ with unitary operators $U$ that commutate with the total Hamiltonian. In detail, the thermal operations act on the Hilbert space of a system $S$ having Hamiltonian $H$, and are defined by performing a Stinespring dilation with an environment $E$ having Hamiltonian $H_{E}$ and prepared in the Gibbs state $\gamma^{E}_{\beta}=e^{-\beta H_{E}}/Z_{E}$ at the temperature $T=\beta^{-1}$, with $Z_{E}=\text{Tr}\left\{e^{-\beta H_{E}}\right\}$, and with a global unitary operator $U$ such that $[U,H+H_{E}]=0$. The thermal operations preserve the Gibbs state of the system $S$, $\Lambda_{th}(\gamma_{\beta})=\gamma_{\beta}$, thus the Gibbs state $\gamma_{\beta}=e^{-\beta H}/Z$, with $Z=\text{Tr}\left\{e^{-\beta H}\right\}$, is the free state. It results that not all the CPTP maps that preserve the Gibbs state $\gamma_{\beta}$, forming the set of Gibbs-preserving operations, are thermal operations (see, e.g., Ref. Faist15 ). Furthermore, the nonequilibrium Helmholtz free energy defined as $F_{\beta}(\rho)=TS(\rho||\gamma_{\beta})-T\ln Z$ does not increase under the action of the thermal operations, i.e., $F_{\beta}(\Lambda_{th}(\rho))\leq F_{\beta}(\rho)$, so that it is a monotone. The maximum work extractable is given by the difference $\Delta F_{\beta}(\rho)=F_{\beta}(\rho)-F_{\beta}(\gamma_{\beta})=TS(\rho||\gamma_{\beta})$. We note that in this case a work extraction is microscopically performed through an environment $E$, which initially is in the equilibrium state $\gamma^{E}_{\beta}$. In contrast, in the following we focus on a reversible work extraction, such that the system $S$ is thermally isolated. Thus, in order to extract a nonzero average work, the unitary time-evolution operator must not commutate with the Hamiltonian. We aim to define some free operations that do not increase the maximum work extractable, which is viewed as a resource. In particular, we find that in the definition of these free operations some sort of temperature plays a pivotal role.
@@ II.2 Work extraction
Let us introduce the resources that we aim to investigate from the resource theories perspective. Following Ref. Allahverdyan04 , we consider a finite quantum system $S$ having a Hilbert space $\mathcal{H}$ with dimension $d$. The Hamiltonian is expressed as
@@
with $\gamma_{\rho}=\gamma_{\beta(\rho)}$ and $\beta(\rho)$ solution of a nonlinear equation for $\beta$, which is $S(\rho_{\beta})=S(\rho)$. Here, we simply refer to $F(\rho)$ as free energy, as in Ref. Bera19 . In general, the maximum work extractable is achieved by maximizing Eq. (4) over all the unitary cycles and it is equal to the ergotropy Allahverdyan04 , here denoted with $\mathcal{E}(\rho)$, i.e.,
@@
where $P_{\rho}=\sum_{k}r_{k}|\epsilon_{k}\rangle\langle\epsilon_{k}|$ is a passive state. In particular, the maximum average work $\mathcal{E}(\rho)$ is obtained for a unitary operator $U_{\rho}=\sum_{k}e^{i\phi_{k}}|\epsilon_{k}\rangle\langle r_{k}|$, where $\phi_{k}$ are some arbitrary real phases, so that the final state $\rho^{\prime}=U_{\rho}\rho U_{\rho}^{\dagger}$ is the passive state $P_{\rho}$. In particular, we note that $P_{U\rho U^{\dagger}}=P_{\rho}$ for any unitary operator $U$. In general, the set $\Pi$ of the passive states is formed by all the states $P$ having zero ergotropy, i.e., such that $\mathcal{E}(P)=0$. In other words, the ergotropy $\mathcal{E}(\rho)$ is zero if and only if the state $\rho$ is passive. All the Gibbs states $\gamma_{\beta}$ with $\beta\geq 0$ are passive states, and in general the set $\Pi$ is formed by all the states $P=\sum_{k}p_{k}|\epsilon_{k}\rangle\langle\epsilon_{k}|$ with $p_{k}\geq p_{k+1}$. If the energy spectrum is degenerate, i.e., $\epsilon_{k}\leq\epsilon_{k+1}$, the set $\Pi$ is formed by all the states $P=\sum_{k}p_{k}U|\epsilon_{k}\rangle\langle\epsilon_{k}|U^{\dagger}$ with $p_{k}\geq p_{k+1}$ and $U$ arbitrary unitary operator such that $[U,H]=0$. In particular, one can check that all the results that we are going to derive trivially hold also for $\epsilon_{k}\leq\epsilon_{k+1}$. We focused on $\epsilon_{k}<\epsilon_{k+1}$ only to introduce the passive state $P_{\rho}$, which is well defined even if there is some degeneracy in the energy.
Let us consider an ensemble of $n$ copies of the system $S$, as in Ref. Alicki13 , where any copy has the Hamiltonian $H$ and is prepared in the state $\rho$. The state $\rho$ is completely passive if the state $\rho^{\otimes n}$ is passive with respect to the total Hamiltonian for any $n$. In general, the set $C$ of the completely passive states is formed by all the Gibbs states $\gamma_{\beta}$ with $\beta\geq 0$. This can be understood by noting that $\mathcal{E}(\rho^{\otimes n})\leq nF(\rho)$ for any $n$. Then, by noting that $F(\rho)=0$ if and only if $\rho=\gamma_{\beta}$ with $\beta\geq 0$, we get $\mathcal{E}(\rho^{\otimes n})=0$ for any $n$ if $\rho$ is a Gibbs state. Furthermore, it is possible to perform a global unitary cycle such that Alicki13 ; Francica24
@@
as $n\to\infty$. Then, if $\rho\neq\gamma_{\beta}$ with $\beta\geq 0$, there exists a sufficient large integer $n$ such that $\mathcal{E}(\rho^{\otimes n})>0$. Thus, we deduce that $\mathcal{E}(\rho^{\otimes n})=0$ for any $n$ if and only if $\rho$ is a Gibbs state. In the end, we note that, since $\gamma_{U\rho U^{\dagger}}=\gamma_{\rho}$ for any unitary operator $U$, the free energy and the ergotropy are related by the equation
@@ III Completely passive states as free states
Then, $\mathcal{O}_{cp}\subseteq\mathcal{M}_{cp}=\{\Lambda|\Lambda(C)\subseteq C\}$ and $M_{cp}(\rho)=\min_{\gamma\in C}S(\rho||\gamma)$ is a monotone, since $M_{cp}(\Lambda(\rho))\leq M_{cp}(\rho)$ for all $\Lambda\in\mathcal{M}_{cp}$. In contrast, by defining the temperature functional $T(\gamma)$ such that $T(\gamma_{\beta})=\beta^{-1}$, the free energy $F(\rho)$ can be expressed as Bera19
@@
and we defined the contraction factor
@@
Thus, the set of free operations $\mathcal{O}_{cp}$ is formed by all the CPTP maps $\Lambda$ that satisfy the conditions (F,i) and (F,iii), where (F,iii) is the condition in Eq. (12). We can see that the condition (F,i) does not imply (F,ii) by considering a map $\Lambda_{\beta}$ such that $\Lambda_{\beta}(\gamma)=\gamma_{\beta}$. Similarly to the maps considered in Refs. Faist15 ; Tajima24 , the map $\Lambda_{\beta}$ can be realized as
@@
for all $\gamma\in C$, where $\sigma_{i}=K_{i}\rho K_{i}^{\dagger}/q_{i}$, $q_{i}=\text{Tr}\left\{K^{\dagger}_{i}K_{i}\rho\right\}$, $\gamma_{i}=K_{i}\gamma K_{i}^{\dagger}/\text{Tr}\left\{K^{\dagger}_{i}K_{i}\gamma\right\}$ and $\Lambda(\rho)=\sum_{i}K_{i}\rho K^{\dagger}_{i}=\sum_{i}q_{i}\sigma_{i}$. For instance, Eq. (25) is satisfied if $T(\gamma_{i})\leq T(\gamma)$ for all $i$, since in general $S(\rho||\gamma)\geq\sum_{i}q_{i}S(\sigma_{i}||\gamma_{i})$ (see, e.g., Ref. Vedral98 ).
@@
We note that the set $C$ is not convex for $d>2$, i.e., the convex combination of two Gibbs states is not always a Gibbs state, and the free energy $F(\rho)$ is not convex, e.g., for $\rho=\sum_{i}p_{i}\gamma_{i}\notin C$, we get $F(\rho)>\sum_{i}p_{i}F(\gamma_{i})=0$ if $\gamma_{i}\in C$ for all $i$. Then, a free operation $\Lambda$ that satisfies the strong monotonicity property is such that $K_{i}\gamma K_{i}^{\dagger}/\text{Tr}\left\{K^{\dagger}_{i}K_{i}\gamma\right\}=\Lambda(\gamma)$ for all $i$. Furthermore, if we consider two systems with Hamiltonian $H_{1}$ and $H_{2}$, and the composite system with Hamiltonian $H=H_{1}+H_{2}$, we recall that the free energy is super-additive under tensor products Bera19 , i.e., $F(\rho_{1}\otimes\rho_{2})\geq F(\rho_{1})+F(\rho_{2})$, and it is additive if $\beta(\rho_{1})=\beta(\rho_{2})$, thus it cannot be considered as a measure.
@@ IV Passive states as free states
Let us try to extend the results achieved to the case of free states that are passive states. We consider the set $\Pi$ of the passive states as the set of the free states. The set $\mathcal{O}_{p}$ of the free operations is formed by all the CPTP maps $\Lambda$ that satisfy (E,i) $\Lambda(\Pi)\subseteq\Pi$ and (E,ii) $\mathcal{E}(\Lambda(\rho))\leq\mathcal{E}(\rho)$ for all $\rho$. Similarly to the free energy case, the condition (E,ii) implies (E,i).
Proof.
If $\mathcal{E}(\Lambda(\rho))\leq\mathcal{E}(\rho)$ for all $\rho$, then $\mathcal{E}(\Lambda(P))\leq\mathcal{E}(P)=0$, and thus $\Lambda(P)\in\Pi$ for all $P\in\Pi$. ∎
Then, $\mathcal{O}_{p}\subseteq\mathcal{M}_{p}=\{\Lambda|\Lambda(\Pi)\subseteq\Pi\}$ and $M_{p}(\rho)=\min_{P\in\Pi}S(\rho||P)$ is a monotone, since $M_{p}(\Lambda(\rho))\leq M_{p}(\rho)$ for all $\Lambda\in\mathcal{M}_{p}$. We note that the ergotropy is equal to the free energy, $\mathcal{E}(\rho)=F(\rho)$, when $P_{\rho}$ is a completely passive state, i.e., $P_{\rho}=\gamma_{\rho}$. Then, to generalize the results achieved for the free energy to the case of the passive states as free states, we try to define something like a nonequilibrium temperature for the passive state $P$ such that it is equal to the temperature functional $T(P)$ when $P=\gamma\in C$. Thus, from the monotone $M_{p}(\rho)$, similarly to Eq. (11), we define a functional
@@
where $T(P|\rho)$ is some functional that plays the role of nonequilibrium temperature for the passive state $P$. Of course there are infinite ways to define a functional $T(P)$ of a passive state $P$ such that $T(P)=\beta^{-1}$ when $P=\gamma_{\beta}$. We refer to this functional $T(P)$ as nonequilibrium temperature since it gives the usual temperature of $P$ when $P$ is a Gibbs state. When we perform the minimization in Eq. (29), we can select a certain definition of the nonequilibrium temperature, among all the possible definitions, depending on the state $\rho$, achieving in this way a nonequilibrium temperature $T(P|\rho)$ that depends on $\rho$. The minimum in Eq. (29) can be achieved for a unique passive state $P^{*}$, such that $\Phi(\rho)=T(P^{*}|\rho)S(\rho||P^{*})$. Given a state $\rho$, we can choose the nonequilibrium temperature $T(P|\rho)$ such that $P^{*}=P_{\rho}$ and the functional in Eq. (29) is equal to the ergotropy, i.e., $\Phi(\rho)=\mathcal{E}(\rho)$. In detail, we find that $P^{*}=P_{\rho}$ and
@@
and we defined the contraction factor